OpenStudy (anonymous):

calculate the mass of a cube of copper, each side of which is 15 cm

OpenStudy (anonymous):

so you can determine the volume, its just 15*15*15=3375 m^3

OpenStudy (anonymous):

\[\huge V_{cube}=(side)^3\]

OpenStudy (anonymous):

after that, you need the density of copper. Multiply the Density by the Volume, and you'll have your solution :)

OpenStudy (anonymous):

okay thanks :). this was physics and i was wondering how to actually get the answer, when i was only given one piece of information. So thanks :)

OpenStudy (anonymous):

Anytime @teeny :)

OpenStudy (anonymous):

\[Mass of Cube = \left( side of cube \right)^{3}\] So the Mass of the copper cube would be (15*15*15)cm = 3,375 \[cm ^{3}\]

OpenStudy (unklerhaukus):

\[\rho=\frac mV\]\[\rho_{\text{Cu}}=8.96 \left[\text {g·cm}^{−3}\right]\]